Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

If Jesse flips a coin seven times in a row, probablity that

Expert replies
by gmattesttaker2 » Sun Jun 15, 2014 4:29 pm
Hello,

Can you please help with this:

If Jesse flips a coin seven times in a row, what is the probability that the result will be heads at least five times?

OA: [spoiler]29/128[/spoiler]

When I saw the word at-least 5 times, I used the following approach:

P(Heads at-least 5 times) = 1 - P(No heads at least 5 times)
= 1 - (1/2.1/2.1/2.1/2.1/2)
= 1 - (1/2)^5
= 1 - 1/32
= 31/32

However, it doesn't seem correct. Can you please assist?

Thanks a lot,
Sri
Join the discussion
Source: — Problem Solving |

by kvcpk » Sun Jun 15, 2014 5:57 pm
gmattesttaker2 wrote:Hello,

Can you please help with this:

If Jesse flips a coin seven times in a row, what is the probability that the result will be heads at least five times?

OA: [spoiler]29/128[/spoiler]

When I saw the word at-least 5 times, I used the following approach:

P(Heads at-least 5 times) = 1 - P(No heads at least 5 times)
= 1 - (1/2.1/2.1/2.1/2.1/2)
= 1 - (1/2)^5
= 1 - 1/32
= 31/32

However, it doesn't seem correct. Can you please assist?

Thanks a lot,
Sri
Hi Sri,

Here is an approach you could follow:
We are looking for a head atleast 5 times.
P(Atleast 5 times head) = P(Exactly 5 times head)+P(Exactly 6 times head) + P(Exactly 7 times head)

Let us calculate P(Exactly 5 times head) :
Probability of Getting 5 heads = 1/2 * 1/2* 1/2 * 1/2 * 1/2 = 1/2^5
Probability of Getting 2 tails = 1/2 * 1/2 = 1/4
Out of the 7 occurrences, We need only 5 occurrences of head. Hence we can obtain them in 7C5 ways.
Therefore, probability = 7C5 * 1/2^5 * 1/4 = 21/128

Applying the same logic Let us calculate P(Exactly 6 times head) :
Probability of Getting 6 heads = 1/2 * 1/2* 1/2 * 1/2 * 1/2 * 1/2= 1/2^6
Probability of Getting 1 tails = 1/2
Out of the 7 occurrences, We need only 6 occurrences of head. Hence we can obtain them in 7C6 ways.
Therefore, probability = 7C6 * 1/2^6 * 1/2 = 7/128

P(Exactly 7 times head) = 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 1/2^7

Adding all these, we get 21/128 + 7/128 + 1/128 = [spoiler]29/128[/spoiler]

Hope this helps!!
"Once you start working on something,
don't be afraid of failure and don't abandon it.
People who work sincerely are the happiest."
Chanakya quotes (Indian politician, strategist and writer, 350 BC-275BC)
Join the discussion

by gmattesttaker2 » Tue Jun 17, 2014 11:24 pm
kvcpk wrote:
gmattesttaker2 wrote:Hello,

Can you please help with this:

If Jesse flips a coin seven times in a row, what is the probability that the result will be heads at least five times?

OA: [spoiler]29/128[/spoiler]

When I saw the word at-least 5 times, I used the following approach:

P(Heads at-least 5 times) = 1 - P(No heads at least 5 times)
= 1 - (1/2.1/2.1/2.1/2.1/2)
= 1 - (1/2)^5
= 1 - 1/32
= 31/32

However, it doesn't seem correct. Can you please assist?

Thanks a lot,
Sri
Hi Sri,

Here is an approach you could follow:
We are looking for a head atleast 5 times.
P(Atleast 5 times head) = P(Exactly 5 times head)+P(Exactly 6 times head) + P(Exactly 7 times head)

Let us calculate P(Exactly 5 times head) :
Probability of Getting 5 heads = 1/2 * 1/2* 1/2 * 1/2 * 1/2 = 1/2^5
Probability of Getting 2 tails = 1/2 * 1/2 = 1/4
Out of the 7 occurrences, We need only 5 occurrences of head. Hence we can obtain them in 7C5 ways.
Therefore, probability = 7C5 * 1/2^5 * 1/4 = 21/128

Applying the same logic Let us calculate P(Exactly 6 times head) :
Probability of Getting 6 heads = 1/2 * 1/2* 1/2 * 1/2 * 1/2 * 1/2= 1/2^6
Probability of Getting 1 tails = 1/2
Out of the 7 occurrences, We need only 6 occurrences of head. Hence we can obtain them in 7C6 ways.
Therefore, probability = 7C6 * 1/2^6 * 1/2 = 7/128

P(Exactly 7 times head) = 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 1/2^7

Adding all these, we get 21/128 + 7/128 + 1/128 = [spoiler]29/128[/spoiler]

Hope this helps!!
Hello KVCP,

Thank you very much for your detailed explanation.

Best Regards,
Sri
Join the discussion

by GMATinsight » Wed Jun 18, 2014 3:55 am
Probability = Favorable outcomes / Total outcomes

With 7 coins tossed, Total Outcomes = 2^7 = 128

Favorable out comes are as follows

1) Exactly 5 heads out of 7 coins = 7C5 = 21
2) Exactly 6 heads out of 7 coins = 7C6 = 7
3) Exactly 7 heads out of 7 coins = 7C7 = 1

Probability = (21+7+1)/128 = 29/128 ANSWER
"GMATinsight"Bhoopendra Singh & Sushma Jha
Most Comprehensive and Affordable Video Course 2000+ CONCEPT Videos and Video Solutions
Whatsapp/Mobile: +91-9999687183 l [email protected]
Contact for One-on-One FREE ONLINE DEMO Class Call/e-mail
Most Efficient and affordable One-On-One Private tutoring fee - US$40-50 per hour
Join the discussion

by Jeff@TargetTestPrep » Mon Apr 23, 2018 4:45 pm
gmattesttaker2 wrote:Hello,

Can you please help with this:

If Jesse flips a coin seven times in a row, what is the probability that the result will be heads at least five times?
We have 3 scenarios:

Scenario 1: 5 heads and 2 tails (e.g., H-H-H-H-H-T-T)

P(H-H-H-H-H-T-T) = (1/2)^7 = 1/128

Number of ways to arrange H-H-H-H-H-T-T = 7C5 = 7!/(5! x 2!) = (7 x 6)/2 = 21

Thus, the probability of this scenario is 1/128 x 21 = 21/128.

Scenario 2: 6 heads and 1 tail (e.g., H-H-H-H-H-H-T)

P(H-H-H-H-H-H-T) = (1/2)^7 = 1/128

Number of ways to arrange H-H-H-H-H-H-T = 7C6 = 7!/6! = 7

Thus the probability of this scenario is 1/128 x 7 = 7/128.

Scenario 3: all 7 heads (i.e., H-H-H-H-H-H-H)

P(H-H-H-H-H-H-H) = (1/2)^7 = 1/128

So, P(at least 5 heads) = 21/128 + 7/128 + 1/128 = [spoiler]29/128[/spoiler]

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion